Barbarian Inc., the maker of personal computer games has been contacted by Super Fast Inc., which has developed a new technology for personal computer gamers which provides a new challenge for gamers. The new technology consists of a seat pad for the gamer’s chair which is connected to a control unit which is connected to the personal computer. When the gamer makes an error while playing Raceway 400 he/she receives a “mild” electric shock through the seat pad. The new technology, which has received extremely positive reviews from a panel of gamers, consists of the Raceway 400 game CD, the seat pad, the control unit and cables. In tests, by Barbarian Inc., the new technology has performed flawlessly, is considered safe and Barbarian Inc. is considering licensing Super Fast’s new gamer technology. The following cost information pertains to the new technology:

Super Fast Raceway 400 Software License.....................$150,000

New Equipment To Produce Seat Pads..............................$75,000

Advertising and promotion................................................$150,000

Material & Royalties “Raceway 400” Game (per 100 CDs) .....$500

Material & Labor For Seat Pad (per 10 pads)..........................$150

Control Unit And Cables (per unit) ...............................................$5

Market research estimates that the market for the new Raceway 400 game and equipment is 1,000,000 units at a retail price of $100, which includes the game CD, seat pad, control unit and cables.

Barbarian Inc. traditionally sells through major retailer clubs like Costco. The club’s traditional margin is 20%.

  1. How much revenue does Barbarian Inc. receive for each unit sold?

  2. What is Barbarian Inc.’s contribution margin per unit?

  3. What is the break-even volume in units?

  4. How many units did Barbarian sell if Total Sales Revenue=$1,000,000?

  5. What is the profit if Barbarian’s Total Sales Revenue= $1,000,000?
Answer #1

Based on the question, we have;

Fixed cost
Swper Fast Raceway 400 License 150,000.00
New equipment for Seat Pads     75,000.00
Advertisement and promotion 150,000.00
Total 375,000.00
Variable Cost
Raceway 400 games per 100 CD          500.00
Material & labour for seat pad (per 10 pads)          150.00
Control unit and Cable              5.00

Hence, per unit variable cost is; 5 + 15 + 5 = 25

Now, the club wants 20% margin, hence selling price for the manufacturer is = Retail price * 0.8 = 80 = (Assuming club's margin is on the retail price)

Hence, profit potential from each unit (Contribution Margin) = 80 - 25 = 55

Now, to cover the fixed cost, the break even volume is;

Break Even volume = Fixed cost/contribution margin = 375000/55 = 6818.18 = 6819 units.

Now, Barbarian's selling price is $80. Hence, if sales volume is 1,000,000.00 then Barbarian has sold;

1000000/80 = 12500 units.

Now, Profit = contribution margin * (Total volume - Break even volume) = 55 * (12500 - 6819) = 312,455